OpenStudy (anonymous):

Solve 5^2x - 4(5^x) = 12

OpenStudy (anonymous):

Because x is an exponent we use logarithms to solve. Take the log of both sides. log(5^2x)-log(4(5^x)=log(12) 2xlog5-(log4+log(5^x))=log(12) 2xlog5-log4-log(5^x)=log12 +log4 +log4 2xlog5-xlog5=log12+log4 xlog5=log(12)(4) /log5 /log5 x=log48/log5

OpenStudy (anonymous):

thank a lot bro ,